### v1

# Small Ramsey numbers for books, wheels, and generalizations

Abstract
This brief note gives several new upper and lower bounds on Ramsey numbers for books and wheels, including tight lower bounds establishing the Ramsey numbers $R(B_{3},B_{6})=19$, $R(B_{4},B_{5})=19$, $R(B_{5},B_{6})=23$, and $R(B_{6},B_{7})=27$, and tight upper and lower bounds establishing $R(W_{5},W_{7})=15$. The new upper bound was obtained using the flag algebra method, and the lower bounds come from constructions discovered using tabu search.
We also explore generalized Ramsey numbers using similar methods. Let $GR(r,K_{s},t)$ denote the minimum number of vertices $n$ such that any $r$-edge-coloring of $K_{n}$ has a copy of $K_{s}$ with at most $t$ colors. We establish $GR(3,K_{4},2)=10,GR(4,K_{4},3)=10$, and some additional bounds.
1 Introduction
The most classical problem in Ramsey theory deals with cliques in 2 colors: namely, it asks us to find the Ramsey number $R(K_{s},K_{t})$, the smallest number $n$ guaranteeing that any red-blue edge-coloring of $K_{n}$ must contain a red copy of $K_{s}$ or a blue copy of $K_{t}$. Equivalently, one may ask for the largest $n$ such that there exists a red-blue edge-coloring of $K_{n-1}$ avoiding the “forbidden” subgraphs $K_{s}$ in red and $K_{r}$ in blue. There are many generalizations of this problem, including to larger numbers of colors, and forbidden subgraphs other than cliques. A dynamic survey by Radziszowski [rad_survey] tracks the state of knowledge in this area.
Here, we will focus on Ramsey numbers where the forbidden subgraphs are wheels or books. The wheel $W_{k}$ consists of a cycle on $k-1$ vertices (“spokes”), together with a central vertex adjacent to all vertices of the cycle. The book $B_{k}$ consists of $k+2$ vertices, an edge $uv$ (the “spine”), and $k$ vertices (“pages”) each adjacent only to $u$ and $v$. See Figure 1 for an illustration.
$W_{6}$
$B_{4}$
Figure 1: Wheel $W_{6}$ and book $B_{4}$.
There are a number of results on Ramsey numbers for books and wheels, summarized in [rad_survey]. In particular, the first and third authors proved upper bounds for a variety of small Ramsey numbers, including for books and wheels, using the flag algebra method [LidP]; we establish the upper bounds here using the same method. A very recent paper [RL_bounds] used reinforcement learning to find Ramsey lower bound constructions for graphs including books and wheels. The idea of finding graph constructions with reinforcement learning was pioneered in [wagner], and it is an exciting and promising technique. But it currently seems to struggle to scale to larger problems in comparison with traditional local search methods [deepmind, RL2, RL1]. Indeed, here we improve each of the lower bounds for books and wheels from [RL_bounds] using tabu search.
We also explore a generalization of Ramsey numbers introduced by Erdős and Shelah [MR0389669, MR0681869]. In classical Ramsey problems, one is looking for a monochromatic copy of a graph $H$ in an $r$-edge-coloring of a graph $K_{n}$. The generalization asks for a copy of $H$ with at most $t$ colors for a fixed $t$. We investigate the case where $H$ is a clique. Let $GR(r,K_{s},t)$ denote the minimum number of vertices $n$ such that any $r$-edge-coloring of $K_{n}$ has a copy of $K_{s}$ with at most $t$ colors. Observe
$GR(r,K_{s},1)=R(\underbrace{K_{s},\ldots,K_{s}}_{r}).$
These generalized Ramsey numbers have been studied asymptotically; for some of the work see [bal2024generalizedramseynumberscycles, gomezleos2024newboundsgeneralizedramsey, MR1645678, MR1656546, MR3872308]. The small values of $r,s,t$ for $GR(r,K_{s},t)$ remained unexplored.
2 Results
We prove the following bounds; asterisks indicate bounds that are tight.
Lower (new) Upper (new) Lower (old) Upper (old)
$R(W_{5},W_{7})$ 15* 15* 14 [RL_bounds] 16 [LidP]
$R(W_{5},W_{9})$ 18 17 [rad_survey]
$R(B_{2},B_{8})$ 21 19 [RS78] 22 [b2_uppers]
$R(B_{2},B_{9})$ 22 21 [RS78] 24 [b2_uppers]
$R(B_{2},B_{10})$ 24 23 [RS78] 26 [b2_uppers]
$R(B_{3},B_{6})$ 19* 17 [RL_bounds] 19 [LidP]
$R(B_{3},B_{7})$ 20
$R(B_{4},B_{5})$ 19* 18 [RL_bounds] 19 [RS78]
$R(B_{4},B_{7})$ 22
$R(B_{5},B_{6})$ 23* 23 [RS78]
$R(B_{5},B_{7})$ 25
$R(B_{6},B_{7})$ 27* 27 [RS78]
$R(B_{6},B_{8})$ 28 29 [RS78]
$R(B_{7},B_{8})$ 29 31 [RS78]
$R(B_{8},B_{8})$ 31 33 [RS78]
$GR(3,K_{4},2)$ 10* 10*
$GR(3,K_{5},2)$ 20 25
$GR(3,K_{6},2)$ 32 54
$GR(4,K_{4},2)$ 15 17
$GR(4,K_{4},3)$ 10* 10*
Table 1: New upper and lower bounds on Ramsey numbers.
3 Method
Each of the lower bounds is established by a construction provided in Appendix A. The constructions were found using tabu search and verified independently in SageMath [sagemath]; the code for books and wheels is available at https://github.com/gwen-mckinley/ramsey-books-wheels. The upper bounds were found by semidefinite programming using the flag algebra method, as described in [LidP]. The certificates for the upper bounds are available at http://lidicky.name/pub/gr. Additionally, for $GR(3,K_{4},2)$ and $GR(4,K_{4},2)$, we were able to use brute force to enumerate all possible constructions for the lower bound. There are 263 such constructions for $GR(3,K_{4},2)$ and only one for $GR(4,K_{4},2)$.
3.1 Tabu search details
To prove a Ramsey lower bound $n$, our goal is to construct an edge-coloring of $K_{n-1}$ avoiding forbidden monochromatic subgraphs (or, in the case of the generalized Ramsey numbers, forbidden subgraphs with too few edge colors). For simplicity, we will restrict the discussion here to the traditional 2-color case, but the ideas are exactly the same for the generalized problem.
In this setting, tabu search works roughly as follows: we start with a random 2-edge-coloring of $K_{n-1}$ (equivalently, a random graph on $n-1$ vertices), and count the number of monochromatic substructures we are trying to avoid – this is the “score” of the graph. Then at each step, we re-color one edge, choosing whichever would produce the lowest-scoring graph, while rejecting any choice that would produce a graph already visited. This restriction is enforced by use of a “tabu set,” which may consist either of forbidden moves (edges recently edited) or of whole graphs/colorings already visited. Here we take the second approach; more details and discussion are given below.
There are many variants of tabu search used in practice, but the one implemented here is very simple, and has two main features:
  1. 1.
    We do not restrict the length of the tabu – i.e., we reject any moves that would produce a graph already visited at any point during the search.
  2. 2.
    We allow the search to run indefinitely, and never restart at a random graph. Instead, we run several instances of the search in parallel, starting from independently sampled random graphs.
This approach has two main advantages: first, because this algorithm has no hyperparameters to tune, it is simple and usable “out of the box” for new problems. Second, by allowing the search to run indefinitely, we avoid inadvertently restarting too quickly. Empirically, it seems that the number of steps required to find a successful construction grows quickly with the problem parameters, and it is not immediately obvious how to predict this value. If we “bid low” and restart the search too quickly, we may consistently cut it short before it is able to reach a promising area of the search space. On the other hand, we observed no clear empirical disadvantage in allowing the search to continue [^2] There are a number of open-source hyperparameter tuning tools supporting this kind of experimentation; we used Optuna here.. Indeed, under some conditions, local search algorithms provably achieve optimal speedup via parallelization (as opposed to restarts); this is dependent on the runtime distribution of the particular algorithm. See Section 4.4 and Chapter 6 of [SLS_book] for further discussion and related empirical results.
We now return to the details of the choice of tabu set. Most variants of tabu search used in the literature store only a list of recent moves (e.g., edges recently edited), rather than whole graphs/configurations, due in large part to memory constraints [tabu_chapter, SLS_book]. However, if only moves are stored, it is important to limit the length of the tabu set – otherwise all possible moves will quickly be forbidden. On the other hand, storing a full representation of each graph visited (and checking the result of each possible move against this set) is very computationally expensive. To avoid this issue, we instead store the graphs by hashing. This offers a much lower computational cost, while still allowing us to avoid all graphs previously visited. (It should be noted, however, that hash collisions may cause us to avoid some other graphs as well – we did not observe issues with this in practice, but it may be worth further study.) This general strategy has been seen very occasionally in the literature [hash]; but again, it is more common to store a tabu set of recent moves, in which case the length of the tabu is an important hyperparameter, which must be tuned correctly or dynamically updated for good performance.
3.2 Scoring functions
Also worth noting: an important ingredient in any local search algorithm is an efficient “scoring” function; here, this means counting the number of “forbidden” substructures in a graph. In fact, rather than re-scoring the entire graph at each step, it is generally more efficient to compute the change in score at each step, since most of the graph will be unaffected by editing a single edge.
For books $B_{k}$, both operations are quite efficient, and the cost is basically independent of $k$ for $k\geq 2$. Explicitly, choosing a copy of $B_{k}$ in a graph $G$ reduces to first choosing a “spine” $(u,v)\in E(G)$, then choosing any $k$ vertices from the common neighborhood $N(u)\cap N(v)$ to serve as the “pages”; see Figure 2.
Figure 2: Book $B_{k}$ in a graph
Notice that for the purpose of counting, we do not need to list all such sets of $k$ vertices; we just need to compute the binomial coefficient $\binom{|N(u)\cap N(v)|}{k}$. Given the value $|N(u)\cap N(v)|$, this operation is very efficient (certainly $O(v(G))$), and in practice, all relevant values of the binomial coefficient can simply be stored in a lookup table. Using similar observations, we can also efficiently compute the “score change” for one step in the search (i.e., the change in the number of books from re-coloring an edge), though the details of the implementation are more involved.
For wheels $W_{k}$, we don’t exploit any especially “nice tricks,” and the cost of our scoring and score change functions are $\Omega(n^{k})$ and $\Omega(n^{k-2})$ respectively – i.e., roughly the cost of enumerating over all subgraphs of the relevant size (where $n$ is the number of vertices in the host graph $G$). Note, however, that significant improvement may be possible; counting wheels reduces largely to counting cycles, where surprisingly efficient algorithms are known [cycle_count].
For generalized Ramsey numbers $GR(r,K_{s},t)$, the situation is slightly different; here, we need to count copies of $K_{s}$ with at most $t$ colors (in an $r$-edge-colored complete graph $G$). Because clique counting is a well-studied problem in non-edge-colored graphs, we use the following strategy: (1) reduce to clique counting in a collection of non-edge-colored graphs, then (2) apply the “Pivoter” algorithm for fast clique-counting [pivoting] to each graph in the collection. This algorithm also facilitates counting cliques that contain a particular edge, which is useful for our “score change” computations. (Note: our code implementing this strategy is a rough prototype, so we have not made it publicly available, but it requires relatively minimal adaptations of the code for books and wheels. There is also an open-source implementation of the Pivoter clique-counting algorithm available online [pivoter_code].)
In slightly more detail, we define the scoring function for the generalized Ramsey problem as follows: for each possible set $\mathcal{C}$ of $t$ colors, we construct the (non-edge-colored) graph $G_{\mathcal{C}}$ consisting of edges whose colors in the original coloring $G$ belong to $\mathcal{C}$. Then we take the “score” of $G$ to be the sum of the number of cliques $K_{s}$ in each of these graphs, that is, $\sum_{\mathcal{C}\in\binom{[r]}{t}}(\text{\# copies of $K_{s}$ in $G_{\mathcal{C}}$})$. Notice that if a copy of $K_{s}$ in $G$ uses strictly fewer than $t$ colors, it will be counted multiple times by this scoring function, once for each set $\mathcal{C}$ containing its colors. This could be avoided, for example by using inclusion-exclusion, but at a higher computational cost. Also, it is reasonable and perhaps even desirable that this scoring function more heavily penalizes copies of $K_{s}$ using smaller numbers of colors – that is, copies that are further from satisfying the desired constraint.
3.3 Flag Algebras
The upper bounds were obtained using the flag algebra method. This method was developed by Razborov [MR2371204], and the main idea is to find a certificate for the bound as a sum of squares, which can be done efficiently using semidefinite programming. We used CSDP [csdp] and MOSEK [mosek] solvers in this work. While the solvers produce numerical results, we created tools to change the solution to an exact arithmetic one in SageMath [sagemath]. The flag algebra method is quite general, and it is naturally applicable to problems where one is interested in asymptotic results. For example, the Turán density of small hypergraphs and its variants [l2norm] is a group of problems where flag algebras led to many new improvements.
The Ramsey problems in this paper are not asymptotic, so the method does not apply directly. Instead, we consider a particular blow-up of the Ramsey graph to transfer the problem from a small finite setting into an asymptotic one. This approach was developed by the first and third author [LidP]. In this work we utilize the machinery developed there to obtain more upper bounds.
4 Further directions
4.1 Lower bounds
To extend the methods in this work for finding lower bounds, there are a number of natural directions to explore. Perhaps most obvious would be to implement everything in a lower-level language – all the code for wheels and books is currently written in Python. In particular, the score-change functions dominate the total cost, and speeding them up would almost certainly improve performance; relevant here is an extensive literature on subgraph counting (see [subgraph_survey]), and it may be possible to simply incorporate existing open-source solutions.
One could also experiment with other local search methods. Previous work on Ramsey lower bounds has used many small variations on tabu search and simulated annealing; some of these methods also achieve considerable speedup by incorporating assumptions about the structure of the final construction (e.g., symmetries, cyclic structure, or subgraph counts), for example, [whatsR55, exoo4_6, exoo13]. More broadly, there is a wide literature on local search algorithms [SLS_book], any of which could potentially be applied here. We experimented briefly with a number of possibilities, but did not find any that offered a clear improvement; nevertheless, it is very likely that an appropriate choice could significantly improve performance.
Another small but potentially effective change could be to incorporate weights in the scoring function – that is, to penalize copies of one forbidden subgraph more than another, depending on their respective sizes and/or how recently they have been edited. This approach has often been used for Ramsey lower bounds, but there appears to be no clear theoretical justification for any particular choice of weights, and some works have even arrived at contradictory conclusions [tankers, exoo13]. This question has been studied more extensively in the SAT literature [clauseWeights, clauseWeightsInitial], and it is possible some of that work could be fruitfully adapted here.
Last but not least, one could try to be be more strategic in choosing where to start a local search. One common approach is to build up from smaller graphs: we first find a construction on a smaller number of vertices avoiding “bad” subgraphs, then add a vertex, and run our local search starting from this new graph (as opposed to starting from a random graph) [whatsR55, exoo4_6]. The hope is that by doing this, we direct the search to a promising area of the (very large) search space. We also tried this method, with limited success, but it is likely worth further exploration.
4.2 Upper bounds
In the flag algebras applications, we used the most straightforward version of the approach introduced in [LidP], which generalizes easily. But when calculating an upper bound for a particular Ramsey number, it is also possible to incorporate more constraints. In particular, one could try to include constraints on smaller graphs implied by smaller Ramsey numbers. For example when bounding $R(5,5)$, one may use $R(4,5)=25$ when looking at neighborhoods of a fixed vertex. This approach has been explored by another researcher [volec], but it requires considerably more effort. In a more distant future it may also be possible to calculate larger instances as computers become faster.
Acknowledgments
The second author would like to thank David Hemminger and Larry Wilson for many helpful and stimulating discussions, and for their consistent encouragement. This work used the computing resources at the Center for Computational Mathematics, University of Colorado Denver, including the Alderaan cluster, supported by the National Science Foundation award OAC-2019089.
[…]
Appendix A Constructions
  • $R(W_{5},W_{7})\geq 15$.
    Adjacency matrix of a 14-vertex graph with no $W_{5}$ in color 0 and no $W_{7}$ in color 1:
    [[0,0,1,1,0,0,1,1,1,1,1,0,0,0],
     [0,0,1,0,0,0,1,1,1,0,1,1,0,0],
     [1,1,0,1,1,1,1,0,0,0,0,0,1,0],
     [1,0,1,0,0,1,1,0,0,0,0,1,1,1],
     [0,0,1,0,0,1,1,0,1,0,0,1,1,0],
     [0,0,1,1,1,0,0,0,0,1,0,1,0,1],
     [1,1,1,1,1,0,0,1,0,0,1,0,0,0],
     [1,1,0,0,0,0,1,0,1,1,0,0,0,1],
     [1,1,0,0,1,0,0,1,0,1,1,1,0,0],
     [1,0,0,0,0,1,0,1,1,0,1,0,1,1],
     [1,1,0,0,0,0,1,0,1,1,0,0,0,1],
     [0,1,0,1,1,1,0,0,1,0,0,0,1,1],
     [0,0,1,1,1,0,0,0,0,1,0,1,0,1],
     [0,0,0,1,0,1,0,1,0,1,1,1,1,0]]
  • $R(W_{5},W_{9})\geq 18$.
    Adjacency matrix of a 17-vertex graph with no $W_{5}$ in color 0 and no $W_{9}$ in color 1:
    [[0,0,0,0,0,0,1,1,1,0,0,1,1,1,0,0,1],
     [0,0,1,1,0,1,0,1,0,1,1,0,0,0,0,1,0],
     [0,1,0,0,1,1,0,0,0,1,1,0,1,1,0,1,0],
     [0,1,0,0,1,1,1,0,0,1,1,0,0,1,0,1,0],
     [0,0,1,1,0,1,0,0,1,1,1,1,0,0,0,1,1],
     [0,1,1,1,1,0,0,1,0,1,0,0,0,0,1,1,0],
     [1,0,0,1,0,0,0,1,1,0,0,1,0,1,1,0,1],
     [1,1,0,0,0,1,1,0,0,0,1,1,1,1,1,0,1],
     [1,0,0,0,1,0,1,0,0,1,0,1,1,0,1,1,1],
     [0,1,1,1,1,1,0,0,1,0,1,0,0,0,0,0,0],
     [0,1,1,1,1,0,0,1,0,1,0,0,0,0,1,1,0],
     [1,0,0,0,1,0,1,1,1,0,0,0,1,1,1,0,0],
     [1,0,1,0,0,0,0,1,1,0,0,1,0,1,1,0,1],
     [1,0,1,1,0,0,1,1,0,0,0,1,1,0,1,0,1],
     [0,0,0,0,0,1,1,1,1,0,1,1,1,1,0,0,1],
     [0,1,1,1,1,1,0,0,1,0,1,0,0,0,0,0,0],
     [1,0,0,0,1,0,1,1,1,0,0,0,1,1,1,0,0]]
  • $R(B_{2},B_{8})\geq 21.$
    Adjacency matrix of a 20-vertex graph with no $B_{2}$ in color 0 and no $B_{8}$ in color 1:
    [[0,1,0,1,0,1,0,1,1,0,0,1,1,0,1,1,1,0,0,1],
     [1,0,0,1,1,1,1,0,1,1,1,0,1,0,0,1,1,0,0,1],
     [0,0,0,0,0,0,1,1,0,1,1,1,1,1,1,0,1,1,1,1],
     [1,1,0,0,1,1,1,1,0,0,0,0,1,0,0,1,0,1,1,1],
     [0,1,0,1,0,1,1,0,1,1,1,0,0,1,0,1,0,1,1,0],
     [1,1,0,1,1,0,0,1,1,0,1,1,0,1,0,1,0,1,0,1],
     [0,1,1,1,1,0,0,0,0,1,1,0,1,0,0,0,1,1,1,1],
     [1,0,1,1,0,1,0,0,0,0,0,1,1,1,1,1,0,1,1,1],
     [1,1,0,0,1,1,0,0,0,1,1,1,0,1,1,1,1,0,0,0],
     [0,1,1,0,1,0,1,0,1,0,1,0,1,1,1,1,1,0,1,0],
     [0,1,1,0,1,1,1,0,1,1,0,1,0,1,0,0,1,1,0,1],
     [1,0,1,0,0,1,0,1,1,0,1,0,0,1,1,0,1,1,0,1],
     [1,1,1,1,0,0,1,1,0,1,0,0,0,0,1,1,1,0,1,1],
     [0,0,1,0,1,1,0,1,1,1,1,1,0,0,1,1,0,1,1,0],
     [1,0,1,0,0,0,0,1,1,1,0,1,1,1,0,1,1,0,1,0],
     [1,1,0,1,1,1,0,1,1,1,0,0,1,1,1,0,0,0,1,0],
     [1,1,1,0,0,0,1,0,1,1,1,1,1,0,1,0,0,0,0,1],
     [0,0,1,1,1,1,1,1,0,0,1,1,0,1,0,0,0,0,1,1],
     [0,0,1,1,1,0,1,1,0,1,0,0,1,1,1,1,0,1,0,0],
     [1,1,1,1,0,1,1,1,0,0,1,1,1,0,0,0,1,1,0,0]]
  • $R(B_{2},B_{9})\geq 22.$
    Adjacency matrix of a 21-vertex graph with no $B_{2}$ in color 0 and no $B_{9}$ in color 1:
    [[0,1,0,1,0,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0,1],
     [1,0,0,1,1,0,0,1,1,1,0,0,0,1,1,0,1,1,0,1,1],
     [0,0,0,0,0,1,0,1,0,0,1,1,1,1,0,1,1,1,1,1,1],
     [1,1,0,0,1,1,1,1,1,0,0,1,1,0,1,0,1,1,0,0,0],
     [0,1,0,1,0,0,1,0,1,1,0,0,1,0,1,1,1,1,1,1,0],
     [1,0,1,1,0,0,1,1,0,0,1,1,1,1,1,1,1,0,0,0,0],
     [1,0,0,1,1,1,0,0,1,1,1,1,1,0,1,1,0,0,1,0,0],
     [1,1,1,1,0,1,0,0,1,0,1,1,0,1,0,1,1,1,0,0,1],
     [1,1,0,1,1,0,1,1,0,1,1,1,0,0,0,1,0,1,1,0,1],
     [1,1,0,0,1,0,1,0,1,0,1,0,0,1,1,1,0,0,1,1,1],
     [1,0,1,0,0,1,1,1,1,1,0,1,0,1,0,1,0,0,1,0,1],
     [1,0,1,1,0,1,1,1,1,0,1,0,1,0,0,0,0,1,1,0,1],
     [1,0,1,1,1,1,1,0,0,0,0,1,0,0,1,1,1,1,1,1,1],
     [1,1,1,0,0,1,0,1,0,1,1,0,0,0,1,1,1,0,0,1,1],
     [1,1,0,1,1,1,1,0,0,1,0,0,1,1,0,1,1,0,0,1,0],
     [0,0,1,0,1,1,1,1,1,1,1,0,1,1,1,0,1,0,1,1,0],
     [0,1,1,1,1,1,0,1,0,0,0,0,1,1,1,1,0,1,0,1,0],
     [0,1,1,1,1,0,0,1,1,0,0,1,1,0,0,0,1,0,1,1,1],
     [0,0,1,0,1,0,1,0,1,1,1,1,1,0,0,1,0,1,0,1,1],
     [0,1,1,0,1,0,0,0,0,1,0,0,1,1,1,1,1,1,1,0,1],
     [1,1,1,0,0,0,0,1,1,1,1,1,1,1,0,0,0,1,1,1,0]]
  • $R(B_{2},B_{10})\geq 24.$
    Adjacency matrix of a 23-vertex graph with no $B_{2}$ in color 0 and no $B_{10}$ in color 1:
    [[0,0,1,1,1,1,1,1,0,1,1,1,1,1,1,0,0,1,0,0,1,1,0],
     [0,0,0,0,0,1,0,1,1,1,0,0,1,0,1,1,1,1,1,1,1,1,0],
     [1,0,0,1,1,0,1,0,0,0,1,1,0,1,1,1,1,1,0,0,0,1,1],
     [1,0,1,0,0,1,1,0,0,0,1,1,1,1,0,1,0,0,0,1,1,1,1],
     [1,0,1,0,0,0,1,1,1,1,1,1,0,1,1,0,1,1,1,0,0,0,1],
     [1,1,0,1,0,0,1,1,0,1,0,0,1,1,1,0,0,0,1,1,1,1,0],
     [1,0,1,1,1,1,0,1,0,0,0,1,0,1,1,0,0,0,1,1,0,1,1],
     [1,1,0,0,1,1,1,0,1,1,0,1,0,0,1,0,0,1,1,1,1,1,0],
     [0,1,0,0,1,0,0,1,0,1,1,1,1,0,0,1,1,1,1,1,1,0,1],
     [1,1,0,0,1,1,0,1,1,0,1,0,1,1,1,0,1,1,1,0,1,0,0],
     [1,0,1,1,1,0,0,0,1,1,0,1,1,1,0,1,1,1,0,0,1,0,1],
     [1,0,1,1,1,0,1,1,1,0,1,0,0,0,0,1,0,1,1,1,1,1,1],
     [1,1,0,1,0,1,0,0,1,1,1,0,0,1,0,1,1,0,1,1,1,0,1],
     [1,0,1,1,1,1,1,0,0,1,1,0,1,0,1,0,1,0,1,0,0,0,1],
     [1,1,1,0,1,1,1,1,0,1,0,0,0,1,0,0,1,1,1,0,0,1,0],
     [0,1,1,1,0,0,0,0,1,0,1,1,1,0,0,0,1,1,0,1,1,1,1],
     [0,1,1,0,1,0,0,0,1,1,1,0,1,1,1,1,0,1,1,0,0,1,1],
     [1,1,1,0,1,0,0,1,1,1,1,1,0,0,1,1,1,0,0,0,1,1,0],
     [0,1,0,0,1,1,1,1,1,1,0,1,1,1,1,0,1,0,0,1,0,0,1],
     [0,1,0,1,0,1,1,1,1,0,0,1,1,0,0,1,0,0,1,0,1,1,1],
     [1,1,0,1,0,1,0,1,1,1,1,1,1,0,0,1,0,1,0,1,0,1,0],
     [1,1,1,1,0,1,1,1,0,0,0,1,0,0,1,1,1,1,0,1,1,0,0],
     [0,0,1,1,1,0,1,0,1,0,1,1,1,1,0,1,1,0,1,1,0,0,0]]
  • $R(B_{3},B_{6})\geq 19$.
    Adjacency matrix of an 18-vertex graph with no $B_{3}$ in color 0 and no $B_{6}$ in color 1:
    [[0,1,1,1,1,0,0,1,1,0,1,1,0,0,0,0,1,1],
     [1,0,0,1,0,1,1,0,0,1,1,1,0,0,1,0,1,1],
     [1,0,0,1,0,0,1,1,1,1,1,0,0,1,0,1,0,1],
     [1,1,1,0,0,1,0,0,1,1,1,0,1,1,0,0,1,0],
     [1,0,0,0,0,1,0,1,1,1,0,1,1,0,0,1,1,1],
     [0,1,0,1,1,0,1,0,1,1,1,1,1,0,0,1,0,0],
     [0,1,1,0,0,1,0,0,1,1,0,1,0,1,1,1,0,1],
     [1,0,1,0,1,0,0,0,0,1,1,1,1,1,1,0,0,1],
     [1,0,1,1,1,1,1,0,0,0,0,1,0,1,0,1,1,0],
     [0,1,1,1,1,1,1,1,0,0,0,0,1,1,0,0,0,1],
     [1,1,1,1,0,1,0,1,0,0,0,1,1,0,1,1,0,0],
     [1,1,0,0,1,1,1,1,1,0,1,0,0,1,1,0,0,0],
     [0,0,0,1,1,1,0,1,0,1,1,0,0,1,1,1,1,0],
     [0,0,1,1,0,0,1,1,1,1,0,1,1,0,1,0,1,0],
     [0,1,0,0,0,0,1,1,0,0,1,1,1,1,0,1,1,1],
     [0,0,1,0,1,1,1,0,1,0,1,0,1,0,1,0,1,1],
     [1,1,0,1,1,0,0,0,1,0,0,0,1,1,1,1,0,1],
     [1,1,1,0,1,0,1,1,0,1,0,0,0,0,1,1,1,0]]
  • $R(B_{3},B_{7})\geq 20$.
    Adjacency matrix of a 19-vertex graph with no $B_{3}$ in color 0 and no $B_{7}$ in color 1:
    [[0,1,1,0,0,0,1,1,1,1,1,0,1,0,1,1,1,0,0],
     [1,0,1,0,1,0,1,1,1,1,0,1,0,0,0,0,1,1,1],
     [1,1,0,1,1,0,0,1,1,0,1,1,0,0,0,1,1,0,0],
     [0,0,1,0,0,1,1,0,1,0,1,1,1,1,0,0,1,1,0],
     [0,1,1,0,0,1,0,1,1,0,1,1,0,0,1,1,0,1,1],
     [0,0,0,1,1,0,1,1,1,1,1,0,0,1,1,0,1,1,0],
     [1,1,0,1,0,1,0,0,0,1,1,0,1,0,1,0,1,1,1],
     [1,1,1,0,1,1,0,0,0,1,1,0,0,1,0,1,1,1,0],
     [1,1,1,1,1,1,0,0,0,0,0,1,1,1,1,0,1,0,0],
     [1,1,0,0,0,1,1,1,0,0,0,0,0,1,1,1,1,0,1],
     [1,0,1,1,1,1,1,1,0,0,0,0,1,0,1,1,0,1,0],
     [0,1,1,1,1,0,0,0,1,0,0,0,1,1,0,1,0,1,1],
     [1,0,0,1,0,0,1,0,1,0,1,1,0,1,1,1,0,1,1],
     [0,0,0,1,0,1,0,1,1,1,0,1,1,0,0,1,1,1,1],
     [1,0,0,0,1,1,1,0,1,1,1,0,1,0,0,1,0,0,1],
     [1,0,1,0,1,0,0,1,0,1,1,1,1,1,1,0,0,0,1],
     [1,1,1,1,0,1,1,1,1,1,0,0,0,1,0,0,0,0,0],
     [0,1,0,1,1,1,1,1,0,0,1,1,1,1,0,0,0,0,1],
     [0,1,0,0,1,0,1,0,0,1,0,1,1,1,1,1,0,1,0]]
  • $R(B_{4},B_{5})\geq 19$.
    Adjacency matrix of an 18-vertex graph with no $B_{4}$ in color 0 and no $B_{5}$ in color 1:
    [[0,1,0,1,1,0,1,1,1,1,1,0,1,0,0,0,1,0],
     [1,0,0,0,1,1,1,0,0,1,0,0,0,1,1,0,1,1],
     [0,0,0,1,0,1,1,1,0,0,1,0,1,1,0,1,1,0],
     [1,0,1,0,0,0,1,1,1,0,0,1,0,1,1,0,1,0],
     [1,1,0,0,0,1,0,1,1,0,0,0,1,1,0,1,0,1],
     [0,1,1,0,1,0,0,0,1,1,1,1,0,1,0,1,1,0],
     [1,1,1,1,0,0,0,0,0,1,0,1,1,1,0,1,0,0],
     [1,0,1,1,1,0,0,0,0,1,1,0,0,1,1,1,0,1],
     [1,0,0,1,1,1,0,0,0,0,0,1,1,0,1,1,1,0],
     [1,1,0,0,0,1,1,1,0,0,1,1,0,0,1,1,0,0],
     [1,0,1,0,0,1,0,1,0,1,0,1,1,0,0,0,1,1],
     [0,0,0,1,0,1,1,0,1,1,1,0,1,1,1,0,0,1],
     [1,0,1,0,1,0,1,0,1,0,1,1,0,0,0,1,0,1],
     [0,1,1,1,1,1,1,1,0,0,0,1,0,0,0,0,0,1],
     [0,1,0,1,0,0,0,1,1,1,0,1,0,0,0,1,1,1],
     [0,0,1,0,1,1,1,1,1,1,0,0,1,0,1,0,0,0],
     [1,1,1,1,0,1,0,0,1,0,1,0,0,0,1,0,0,1],
     [0,1,0,0,1,0,0,1,0,0,1,1,1,1,1,0,1,0]]
  • $R(B_{4},B_{7})\geq 22$.
    Adjacency matrix of a 21-vertex graph with no $B_{4}$ in color 0 and no $B_{7}$ in color 1:
    [[0,0,0,1,0,0,1,0,0,0,1,1,1,1,1,1,1,0,1,0,0],
     [0,0,1,0,1,1,1,1,1,0,0,1,1,0,0,1,0,1,1,0,1],
     [0,1,0,1,0,0,1,1,0,1,1,1,0,0,0,1,0,1,0,1,1],
     [1,0,1,0,0,0,1,1,1,1,1,0,0,0,0,1,1,1,1,0,0],
     [0,1,0,0,0,1,1,0,1,1,0,0,1,0,1,0,1,0,1,1,1],
     [0,1,0,0,1,0,0,1,1,1,0,1,1,1,0,1,1,0,0,1,0],
     [1,1,1,1,1,0,0,0,0,0,0,1,0,0,1,1,1,0,1,1,1],
     [0,1,1,1,0,1,0,0,1,0,1,1,0,0,0,0,1,1,1,1,0],
     [0,1,0,1,1,1,0,1,0,1,0,1,0,1,1,0,0,1,1,0,0],
     [0,0,1,1,1,1,0,0,1,0,0,0,0,1,1,1,1,1,0,1,1],
     [1,0,1,1,0,0,0,1,0,0,0,1,1,1,1,0,0,1,0,1,1],
     [1,1,1,0,0,1,1,1,1,0,1,0,0,1,1,1,0,0,0,1,0],
     [1,1,0,0,1,1,0,0,0,0,1,0,0,1,0,1,1,1,1,0,1],
     [1,0,0,0,0,1,0,0,1,1,1,1,1,0,1,1,0,1,0,0,1],
     [1,0,0,0,1,0,1,0,1,1,1,1,0,1,0,0,0,0,1,1,1],
     [1,1,1,1,0,1,1,0,0,1,0,1,1,1,0,0,1,0,0,0,0],
     [1,0,0,1,1,1,1,1,0,1,0,0,1,0,0,1,0,0,1,1,0],
     [0,1,1,1,0,0,0,1,1,1,1,0,1,1,0,0,0,0,1,0,1],
     [1,1,0,1,1,0,1,1,1,0,0,0,1,0,1,0,1,1,0,0,0],
     [0,0,1,0,1,1,1,1,0,1,1,1,0,0,1,0,1,0,0,0,1],
     [0,1,1,0,1,0,1,0,0,1,1,0,1,1,1,0,0,1,0,1,0]]
  • $R(B_{5},B_{6})\geq 23$.
    Adjacency matrix of a 22-vertex graph with no $B_{5}$ in color 0 and no $B_{6}$ in color 1:
    [[0,0,1,1,1,0,1,0,1,0,1,0,0,1,1,1,0,0,1,0,1,0],
     [0,0,1,1,0,1,0,1,1,1,1,0,1,0,0,1,0,1,0,0,0,1],
     [1,1,0,1,0,0,1,1,1,0,1,0,1,1,0,0,1,0,0,1,0,0],
     [1,1,1,0,1,1,0,0,1,0,0,0,0,0,1,1,1,1,0,1,0,0],
     [1,0,0,1,0,1,0,1,0,0,0,1,0,1,1,1,1,0,0,0,1,1],
     [0,1,0,1,1,0,0,0,1,1,1,1,0,1,0,0,1,0,1,0,0,1],
     [1,0,1,0,0,0,0,1,1,1,0,0,0,1,1,0,1,1,1,0,0,1],
     [0,1,1,0,1,0,1,0,0,0,0,1,1,1,0,1,1,1,0,0,0,1],
     [1,1,1,1,0,1,1,0,0,1,0,1,1,0,1,0,0,0,1,0,0,0],
     [0,1,0,0,0,1,1,0,1,0,0,0,1,1,1,1,0,0,0,1,1,1],
     [1,1,1,0,0,1,0,0,0,0,0,1,0,1,0,0,0,1,1,1,1,1],
     [0,0,0,0,1,1,0,1,1,0,1,0,1,1,1,0,0,1,1,1,0,0],
     [0,1,1,0,0,0,0,1,1,1,0,1,0,0,0,1,1,0,1,1,1,0],
     [1,0,1,0,1,1,1,1,0,1,1,1,0,0,0,1,0,0,0,1,0,0],
     [1,0,0,1,1,0,1,0,1,1,0,1,0,0,0,0,0,1,0,1,1,1],
     [1,1,0,1,1,0,0,1,0,1,0,0,1,1,0,0,0,1,1,0,1,0],
     [0,0,1,1,1,1,1,1,0,0,0,0,1,0,0,0,0,0,1,1,1,1],
     [0,1,0,1,0,0,1,1,0,0,1,1,0,0,1,1,0,0,1,1,0,1],
     [1,0,0,0,0,1,1,0,1,0,1,1,1,0,0,1,1,1,0,0,1,0],
     [0,0,1,1,0,0,0,0,0,1,1,1,1,1,1,0,1,1,0,0,1,0],
     [1,0,0,0,1,0,0,0,0,1,1,0,1,0,1,1,1,0,1,1,0,1],
     [0,1,0,0,1,1,1,1,0,1,1,0,0,0,1,0,1,1,0,0,1,0]]
  • $R(B_{5},B_{7})\geq 25$.
    Adjacency matrix of a 24-vertex graph with no $B_{5}$ in color 0 and no $B_{7}$ in color 1:
    [[0,0,0,0,0,1,0,1,1,1,0,0,1,1,1,1,0,0,1,1,1,0,1,1],
     [0,0,0,1,0,0,0,0,1,1,1,0,0,1,1,0,1,1,0,1,0,1,1,1],
     [0,0,0,0,1,0,1,1,1,1,1,0,1,0,0,1,1,0,0,1,1,1,0,0],
     [0,1,0,0,0,1,0,0,0,1,1,1,1,0,0,1,0,1,1,0,1,1,1,0],
     [0,0,1,0,0,1,1,1,0,1,0,1,0,1,1,1,1,1,0,0,0,0,1,0],
     [1,0,0,1,1,0,0,1,1,0,0,1,1,1,1,0,0,1,0,0,1,1,0,0],
     [0,0,1,0,1,0,0,0,0,0,1,1,0,1,1,1,0,1,1,1,1,1,0,1],
     [1,0,1,0,1,1,0,0,1,1,0,1,0,0,0,0,1,0,0,0,1,1,1,1],
     [1,1,1,0,0,1,0,1,0,0,1,0,1,0,1,1,1,1,0,0,0,1,0,1],
     [1,1,1,1,1,0,0,1,0,0,1,0,0,1,0,1,0,0,0,0,1,0,1,1],
     [0,1,1,1,0,0,1,0,1,1,0,1,1,1,0,1,0,0,0,0,0,1,0,1],
     [0,0,0,1,1,1,1,1,0,0,1,0,1,1,0,0,1,0,1,0,0,1,1,1],
     [1,0,1,1,0,1,0,0,1,0,1,1,0,1,0,1,1,0,1,1,0,0,0,0],
     [1,1,0,0,1,1,1,0,0,1,1,1,1,0,1,0,0,0,0,1,0,0,0,1],
     [1,1,0,0,1,1,1,0,1,0,0,0,0,1,0,1,0,1,0,1,0,1,1,0],
     [1,0,1,1,1,0,1,0,1,1,1,0,1,0,1,0,0,1,1,0,0,0,1,0],
     [0,1,1,0,1,0,0,1,1,0,0,1,1,0,0,0,0,1,1,1,0,0,1,1],
     [0,1,0,1,1,1,1,0,1,0,0,0,0,0,1,1,1,0,1,0,1,0,0,1],
     [1,0,0,1,0,0,1,0,0,0,0,1,1,0,0,1,1,1,0,1,1,0,1,1],
     [1,1,1,0,0,0,1,0,0,0,0,0,1,1,1,0,1,0,1,0,1,1,1,0],
     [1,0,1,1,0,1,1,1,0,1,0,0,0,0,0,0,0,1,1,1,0,1,0,1],
     [0,1,1,1,0,1,1,1,1,0,1,1,0,0,1,0,0,0,0,1,1,0,1,0],
     [1,1,0,1,1,0,0,1,0,1,0,1,0,0,1,1,1,0,1,1,0,1,0,0],
     [1,1,0,0,0,0,1,1,1,1,1,1,0,1,0,0,1,1,1,0,1,0,0,0]]
  • $R(B_{6},B_{7})\geq 27$.
    Adjacency matrix of a 26-vertex graph with no $B_{6}$ in color 0 and no $B_{7}$ in color 1:
    [[0,1,1,0,1,0,0,1,1,0,0,1,0,1,0,1,0,0,1,1,0,1,0,0,1,1],
     [1,0,0,0,0,1,1,1,0,1,0,1,1,1,1,1,0,0,0,1,0,0,0,1,1,0],
     [1,0,0,1,1,0,1,1,0,0,1,1,1,0,0,0,0,0,1,1,1,0,1,0,1,0],
     [0,0,1,0,1,1,0,0,1,1,1,0,1,0,1,0,0,1,1,1,0,0,0,1,1,0],
     [1,0,1,1,0,1,0,1,1,1,1,0,0,1,1,1,0,0,0,0,1,0,0,0,0,1],
     [0,1,0,1,1,0,0,0,1,1,1,1,1,1,0,0,0,0,0,0,1,1,1,1,0,0],
     [0,1,1,0,0,0,0,0,1,1,1,1,1,0,1,1,1,0,0,1,0,0,1,0,0,1],
     [1,1,1,0,1,0,0,0,0,1,0,0,0,0,1,1,1,0,1,0,1,0,1,1,1,0],
     [1,0,0,1,1,1,1,0,0,1,0,0,0,0,1,0,1,0,0,1,0,1,1,0,1,1],
     [0,1,0,1,1,1,1,1,1,0,0,1,1,0,0,1,1,1,1,0,0,0,0,0,0,0],
     [0,0,1,1,1,1,1,0,0,0,0,1,0,1,0,1,0,1,0,0,0,0,1,1,1,1],
     [1,1,1,0,0,1,1,0,0,1,1,0,0,1,1,0,0,1,1,0,0,1,1,0,0,0],
     [0,1,1,1,0,1,1,0,0,1,0,0,0,1,0,0,1,0,1,1,1,0,0,1,0,1],
     [1,1,0,0,1,1,0,0,0,0,1,1,1,0,1,0,1,1,0,0,1,0,0,0,1,1],
     [0,1,0,1,1,0,1,1,1,0,0,1,0,1,0,0,0,1,1,1,1,0,1,1,0,1],
     [1,1,0,0,1,0,1,1,0,1,1,0,0,0,0,0,0,1,0,1,1,1,0,1,0,1],
     [0,0,0,0,0,0,1,1,1,1,0,0,1,1,0,0,0,1,1,0,1,1,1,0,1,1],
     [0,0,0,1,0,0,0,0,0,1,1,1,0,1,1,1,1,0,1,1,1,1,0,0,1,0],
     [1,0,1,1,0,0,0,1,0,1,0,1,1,0,1,0,1,1,0,0,0,1,0,1,0,1],
     [1,1,1,1,0,0,1,0,1,0,0,0,1,0,1,1,0,1,0,0,1,1,0,0,1,0],
     [0,0,1,0,1,1,0,1,0,0,0,0,1,1,1,1,1,1,0,1,0,1,1,0,0,0],
     [1,0,0,0,0,1,0,0,1,0,0,1,0,0,0,1,1,1,1,1,1,0,1,1,0,1],
     [0,0,1,0,0,1,1,1,1,0,1,1,0,0,1,0,1,0,0,0,1,1,0,1,1,0],
     [0,1,0,1,0,1,0,1,0,0,1,0,1,0,1,1,0,0,1,0,0,1,1,0,1,1],
     [1,1,1,1,0,0,0,1,1,0,1,0,0,1,0,0,1,1,0,1,0,0,1,1,0,0],
     [1,0,0,0,1,0,1,0,1,0,1,0,1,1,1,1,1,0,1,0,0,1,0,1,0,0]]
  • $R(B_{6},B_{8})\geq 28$.
    Adjacency matrix of a 27-vertex graph with no $B_{6}$ in color 0 and no $B_{8}$ in color 1:
    [[0,1,0,1,0,1,0,0,1,0,1,0,1,1,0,1,1,1,0,0,0,0,1,1,0,1,1],
     [1,0,1,1,0,0,0,1,1,0,1,0,1,0,1,1,1,0,1,1,1,0,0,0,1,1,0],
     [0,1,0,1,1,0,0,0,1,0,0,1,0,0,1,0,1,0,1,1,0,1,1,1,0,0,1],
     [1,1,1,0,1,0,1,0,0,1,0,0,1,1,1,1,0,0,0,0,0,1,1,0,1,1,0],
     [0,0,1,1,0,0,0,1,0,0,1,0,0,1,1,0,0,1,1,1,1,1,1,0,0,1,1],
     [1,0,0,0,0,0,1,1,1,1,1,0,1,1,0,0,0,0,1,1,0,0,1,0,1,0,1],
     [0,0,0,1,0,1,0,1,1,1,1,1,0,1,1,1,0,0,1,0,0,0,0,1,1,1,0],
     [0,1,0,0,1,1,1,0,1,0,0,1,1,1,1,0,1,1,0,1,0,0,0,0,1,1,0],
     [1,1,1,0,0,1,1,1,0,0,0,1,0,0,0,0,0,0,1,1,0,0,1,1,0,1,1],
     [0,0,0,1,0,1,1,0,0,0,1,1,1,0,0,0,1,0,1,0,1,1,1,0,1,1,0],
     [1,1,0,0,1,1,1,0,0,1,0,0,0,1,0,1,1,1,1,1,1,0,1,0,0,0,0],
     [0,0,1,0,0,0,1,1,1,1,0,0,0,0,0,1,1,1,0,1,1,1,1,1,1,0,0],
     [1,1,0,1,0,1,0,1,0,1,0,0,0,1,0,0,1,0,0,0,1,1,0,0,1,1,1],
     [1,0,0,1,1,1,1,1,0,0,1,0,1,0,1,1,1,0,0,1,0,1,0,0,0,0,1],
     [0,1,1,1,1,0,1,1,0,0,0,0,0,1,0,1,1,1,1,0,0,0,0,1,1,0,1],
     [1,1,0,1,0,0,1,0,0,0,1,1,0,1,1,0,0,1,0,1,1,1,0,1,1,0,0],
     [1,1,1,0,0,0,0,1,0,1,1,1,1,1,1,0,0,1,1,0,0,1,1,1,0,0,0],
     [1,0,0,0,1,0,0,1,0,0,1,1,0,0,1,1,1,0,0,0,1,0,1,1,1,1,1],
     [0,1,1,0,1,1,1,0,1,1,1,0,0,0,1,0,1,0,0,0,1,0,0,1,0,1,1],
     [0,1,1,0,1,1,0,1,1,0,1,1,0,1,0,1,0,0,0,0,1,1,1,0,0,0,0],
     [0,1,0,0,1,0,0,0,0,1,1,1,1,0,0,1,0,1,1,1,0,1,0,0,1,1,1],
     [0,0,1,1,1,0,0,0,0,1,0,1,1,1,0,1,1,0,0,1,1,0,0,1,0,1,0],
     [1,0,1,1,1,1,0,0,1,1,1,1,0,0,0,0,1,1,0,1,0,0,0,0,1,0,1],
     [1,0,1,0,0,0,1,0,1,0,0,1,0,0,1,1,1,1,1,0,0,1,0,0,0,1,1],
     [0,1,0,1,0,1,1,1,0,1,0,1,1,0,1,1,0,1,0,0,1,0,1,0,0,0,1],
     [1,1,0,1,1,0,1,1,1,1,0,0,1,0,0,0,0,1,1,0,1,1,0,1,0,0,0],
     [1,0,1,0,1,1,0,0,1,0,0,0,1,1,1,0,0,1,1,0,1,0,1,1,1,0,0]]
  • $R(B_{7},B_{8})\geq 29$.
    Adjacency matrix of a 28-vertex graph with no $B_{7}$ in color 0 and no $B_{8}$ in color 1:
    [[0,0,0,1,0,0,1,1,1,0,0,0,1,0,1,0,1,1,0,1,1,0,1,1,0,1,1,1],
     [0,0,0,1,1,0,1,1,1,0,1,1,0,1,0,1,0,1,0,0,1,0,1,1,1,0,0,0],
     [0,0,0,0,1,0,0,0,0,1,0,1,1,1,0,1,1,1,1,1,0,1,0,1,0,1,1,0],
     [1,1,0,0,0,1,1,0,1,1,0,1,1,0,1,0,0,1,0,0,0,1,1,0,1,1,0,0],
     [0,1,1,0,0,1,1,1,0,0,0,1,1,0,1,1,0,1,1,0,0,0,0,1,1,1,0,1],
     [0,0,0,1,1,0,0,0,1,1,1,0,1,1,1,0,0,1,0,1,0,0,0,1,1,1,0,1],
     [1,1,0,1,1,0,0,0,0,1,0,1,0,1,1,0,1,0,1,0,1,0,1,1,0,1,0,0],
     [1,1,0,0,1,0,0,0,1,1,0,0,0,0,1,1,0,1,1,0,0,1,1,1,0,0,1,1],
     [1,1,0,1,0,1,0,1,0,0,0,0,1,1,1,1,1,0,1,1,0,0,0,1,1,0,0,0],
     [0,0,1,1,0,1,1,1,0,0,0,1,1,1,1,0,1,0,0,0,0,1,1,1,0,0,1,0],
     [0,1,0,0,0,1,0,0,0,0,0,0,1,1,1,1,0,0,1,0,1,1,1,0,1,1,1,1],
     [0,1,1,1,1,0,1,0,0,1,0,0,1,0,0,1,1,1,0,1,0,0,1,0,1,0,0,0],
     [1,0,1,1,1,1,0,0,1,1,1,1,0,1,0,1,0,0,1,0,0,0,1,0,0,0,0,1],
     [0,1,1,0,0,1,1,0,1,1,1,0,1,0,0,0,1,1,1,0,1,0,0,1,0,0,1,0],
     [1,0,0,1,1,1,1,1,1,1,1,0,0,0,0,1,1,0,0,0,1,1,0,0,1,0,1,0],
     [0,1,1,0,1,0,0,1,1,0,1,1,1,0,1,0,0,0,0,1,1,1,0,1,0,0,0,0],
     [1,0,1,0,0,0,1,0,1,1,0,1,0,1,1,0,0,1,1,1,1,0,0,0,1,0,0,1],
     [1,1,1,1,1,1,0,1,0,0,0,1,0,1,0,0,1,0,0,0,1,1,0,0,1,0,0,1],
     [0,0,1,0,1,0,1,1,1,0,1,0,1,1,0,0,1,0,0,0,0,1,0,0,1,1,1,1],
     [1,0,1,0,0,1,0,0,1,0,0,1,0,0,0,1,1,0,0,0,1,0,1,1,1,1,1,0],
     [1,1,0,0,0,0,1,0,0,0,1,0,0,1,1,1,1,1,0,1,0,1,0,0,0,1,0,1],
     [0,0,1,1,0,0,0,1,0,1,1,0,0,0,1,1,0,1,1,0,1,0,0,0,0,1,1,1],
     [1,1,0,1,0,0,1,1,0,1,1,1,1,0,0,0,0,0,0,1,0,0,0,0,1,0,1,1],
     [1,1,1,0,1,1,1,1,1,1,0,0,0,1,0,1,0,0,0,1,0,0,0,0,0,1,1,0],
     [0,1,0,1,1,1,0,0,1,0,1,1,0,0,1,0,1,1,1,1,0,0,1,0,0,0,1,0],
     [1,0,1,1,1,1,1,0,0,0,1,0,0,0,0,0,0,0,1,1,1,1,0,1,0,0,1,1],
     [1,0,1,0,0,0,0,1,0,1,1,0,0,1,1,0,0,0,1,1,0,1,1,1,1,1,0,0],
     [1,0,0,0,1,1,0,1,0,0,1,0,1,0,0,0,1,1,1,0,1,1,1,0,0,1,0,0]]
  • $R(B_{8},B_{8})\geq 31$.
    Adjacency matrix of a 30-vertex graph with no $B_{8}$ in color 0 and no $B_{8}$ in color 1:
    [[0,0,1,0,0,0,1,1,1,1,0,0,0,0,1,1,1,0,1,1,1,1,0,1,0,1,1,0,1,0],
     [0,0,1,0,1,0,0,0,1,1,1,0,1,1,0,1,1,1,0,1,0,0,0,1,0,1,1,1,0,0],
     [1,1,0,0,0,1,1,0,1,0,1,1,0,1,1,1,0,1,0,1,0,1,0,0,0,0,0,0,0,1],
     [0,0,0,0,0,1,1,0,0,0,1,0,1,1,1,0,1,0,1,0,1,1,1,1,0,0,1,1,0,0],
     [0,1,0,0,0,1,0,1,1,1,1,1,0,0,1,0,1,0,0,0,1,0,0,1,1,1,1,0,0,0],
     [0,0,1,1,1,0,0,0,1,0,1,0,1,0,1,0,0,1,1,1,1,1,1,0,1,0,0,0,1,0],
     [1,0,1,1,0,0,0,1,0,0,1,0,0,1,0,0,0,1,1,1,1,1,0,1,0,0,0,1,0,1],
     [1,0,0,0,1,0,1,0,1,1,0,0,1,1,0,0,0,1,0,0,1,0,1,1,0,0,1,0,1,1],
     [1,1,1,0,1,1,0,1,0,1,0,0,1,0,1,0,1,0,0,1,0,0,1,1,0,0,0,0,0,1],
     [1,1,0,0,1,0,0,1,1,0,0,0,1,0,1,1,1,1,1,0,1,0,0,0,1,0,0,1,0,0],
     [0,1,1,1,1,1,1,0,0,0,0,1,0,0,0,1,1,0,0,1,1,0,0,0,0,0,1,1,1,1],
     [0,0,1,0,1,0,0,0,0,0,1,0,0,1,0,1,1,1,1,0,1,0,1,1,1,1,0,0,1,1],
     [0,1,0,1,0,1,0,1,1,1,0,0,0,1,0,0,1,0,1,0,0,1,1,0,1,0,1,0,1,1],
     [0,1,1,1,0,0,1,1,0,0,0,1,1,0,1,1,0,1,1,0,0,0,1,1,0,0,1,0,0,0],
     [1,0,1,1,1,1,0,0,1,1,0,0,0,1,0,1,0,0,0,0,0,1,1,0,1,1,1,1,0,0],
     [1,1,1,0,0,0,0,0,0,1,1,1,0,1,1,0,0,0,1,0,0,0,0,0,1,0,1,1,1,1],
     [1,1,0,1,1,0,0,0,1,1,1,1,1,0,0,0,0,0,0,0,1,1,1,0,0,1,0,1,0,1],
     [0,1,1,0,0,1,1,1,0,1,0,1,0,1,0,0,0,0,0,1,1,1,1,0,1,1,0,1,0,0],
     [1,0,0,1,0,1,1,0,0,1,0,1,1,1,0,1,0,0,0,1,1,0,0,1,1,0,0,1,1,0],
     [1,1,1,0,0,1,1,0,1,0,1,0,0,0,0,0,0,1,1,0,0,0,1,1,0,1,0,1,1,0],
     [1,0,0,1,1,1,1,1,0,1,1,1,0,0,0,0,1,1,1,0,0,1,0,0,0,0,0,0,1,0],
     [1,0,1,1,0,1,1,0,0,0,0,0,1,0,1,0,1,1,0,0,1,0,0,0,1,1,1,0,0,1],
     [0,0,0,1,0,1,0,1,1,0,0,1,1,1,1,0,1,1,0,1,0,0,0,0,0,1,0,1,1,0],
     [1,1,0,1,1,0,1,1,1,0,0,1,0,1,0,0,0,0,1,1,0,0,0,0,1,1,1,0,0,1],
     [0,0,0,0,1,1,0,0,0,1,0,1,1,0,1,1,0,1,1,0,0,1,0,1,0,1,0,1,0,1],
     [1,1,0,0,1,0,0,0,0,0,0,1,0,0,1,0,1,1,0,1,0,1,1,1,1,0,1,1,1,0],
     [1,1,0,1,1,0,0,1,0,0,1,0,1,1,1,1,0,0,0,0,0,1,0,1,0,1,0,0,1,0],
     [0,1,0,1,0,0,1,0,0,1,1,0,0,0,1,1,1,1,1,1,0,0,1,0,1,1,0,0,0,1],
     [1,0,0,0,0,1,0,1,0,0,1,1,1,0,0,1,0,0,1,1,1,0,1,0,0,1,1,0,0,1],
     [0,0,1,0,0,0,1,1,1,0,1,1,1,0,0,1,1,0,0,0,0,1,0,1,1,0,0,1,1,0]]
  • $GR(3,K_{4},2)=10$.
    Adjacency matrix of a 3-edge-colored 9-vertex graph with each $K_{4}$ containing more than 2 colors. There are 263 such colorings (up to color permutations).
    [[0,2,0,0,0,1,1,1,2],
     [2,0,1,1,2,1,0,0,2],
     [0,1,0,2,0,1,0,2,1],
     [0,1,2,0,1,0,1,2,0],
     [0,2,0,1,0,2,2,0,1],
     [1,1,1,0,2,0,2,0,0],
     [1,0,0,1,2,2,0,1,0],
     [1,0,2,2,0,0,1,0,1],
     [2,2,1,0,1,0,0,1,0]]
  • $GR(3,K_{5},2)\geq 20$
    Adjacency matrix of a 3-edge-colored 19-vertex graph with each $K_{5}$ containing more than 2 colors.
    [[0,0,1,1,0,1,1,2,0,2,1,1,2,2,0,2,0,2,2],
     [0,0,0,2,0,2,1,1,2,2,0,1,1,1,1,2,0,0,1],
     [1,0,0,0,2,1,2,2,1,1,2,1,0,0,0,0,1,0,2],
     [1,2,0,0,1,2,0,0,2,2,1,0,1,0,1,2,0,0,1],
     [0,0,2,1,0,1,2,2,0,1,2,0,1,1,0,1,0,1,2],
     [1,2,1,2,1,0,0,2,0,0,2,1,0,2,2,1,2,1,0],
     [1,1,2,0,2,0,0,1,0,2,2,1,1,1,2,0,2,0,0],
     [2,1,2,0,2,2,1,0,1,0,0,0,1,2,0,1,1,0,2],
     [0,2,1,2,0,0,0,1,0,1,1,2,2,1,2,0,1,2,0],
     [2,2,1,2,1,0,2,0,1,0,0,0,0,1,0,2,1,1,0],
     [1,0,2,1,2,2,2,0,1,0,0,0,2,0,1,0,2,2,1],
     [1,1,1,0,0,1,1,0,2,0,0,0,2,0,2,0,2,2,1],
     [2,1,0,1,1,0,1,1,2,0,2,2,0,0,0,1,1,2,0],
     [2,1,0,0,1,2,1,2,1,1,0,0,0,0,2,0,2,1,2],
     [0,1,0,1,0,2,2,0,2,0,1,2,0,2,0,0,2,1,1],
     [2,2,0,2,1,1,0,1,0,2,0,0,1,0,0,0,1,1,2],
     [0,0,1,0,0,2,2,1,1,1,2,2,1,2,2,1,0,0,1],
     [2,0,0,0,1,1,0,0,2,1,2,2,2,1,1,1,0,0,1],
     [2,1,2,1,2,0,0,2,0,0,1,1,0,2,1,2,1,1,0]]
  • $GR(3,K_{6},2)\geq 32$
    Adjacency matrix of a 3-edge-colored 31-vertex graph with each $K_{6}$ containing more than 2 colors.
    [[0,0,0,0,0,2,1,2,1,1,1,1,0,1,0,2,2,2,0,0,1,0,2,1,2,2,0,0,2,1,1],
     [0,0,2,1,0,2,0,2,2,1,2,2,2,1,1,1,2,2,2,2,1,0,0,0,1,1,1,1,0,0,2],
     [0,2,0,2,1,1,0,0,1,0,2,2,0,0,2,2,1,1,0,1,0,2,1,0,0,1,1,1,1,2,2],
     [0,1,2,0,1,0,1,0,0,1,1,2,0,1,2,0,1,1,1,0,1,2,2,0,2,2,0,0,0,2,1],
     [0,0,1,1,0,1,1,2,1,1,2,1,1,2,0,2,1,2,0,2,2,0,2,2,1,2,1,0,0,0,2],
     [2,2,1,0,1,0,1,2,2,0,1,0,2,2,0,0,1,0,2,2,1,0,2,0,1,1,1,0,1,1,0],
     [1,0,0,1,1,1,0,1,1,2,0,1,0,0,2,2,0,1,1,0,2,2,0,1,1,0,2,2,1,0,1],
     [2,2,0,0,2,2,1,0,2,0,2,0,0,0,1,1,2,2,1,2,0,1,1,0,1,1,0,2,0,2,1],
     [1,2,1,0,1,2,1,2,0,2,2,1,2,2,1,0,1,2,2,0,2,1,2,1,0,2,0,1,1,0,0],
     [1,1,0,1,1,0,2,0,2,0,2,2,0,1,0,1,1,1,1,1,2,1,0,0,0,2,2,2,0,2,1],
     [1,2,2,1,2,1,0,2,2,2,0,2,2,2,2,2,0,2,0,0,1,2,0,1,1,0,1,0,2,1,2],
     [1,2,2,2,1,0,1,0,1,2,2,0,1,0,0,0,1,0,1,2,0,1,1,1,1,2,1,2,0,2,0],
     [0,2,0,0,1,2,0,0,2,0,2,1,0,0,1,2,1,1,2,1,0,1,1,0,1,2,1,1,2,2,2],
     [1,1,0,1,2,2,0,0,2,1,2,0,0,0,1,1,1,0,2,2,0,2,2,2,2,1,1,1,0,1,0],
     [0,1,2,2,0,0,2,1,1,0,2,0,1,1,0,1,2,0,0,1,2,2,1,2,0,0,2,1,1,0,0],
     [2,1,2,0,2,0,2,1,0,1,2,0,2,1,1,0,0,0,0,0,1,1,0,0,1,1,0,2,2,0,2],
     [2,2,1,1,1,1,0,2,1,1,0,1,1,1,2,0,0,2,0,2,2,0,0,2,2,0,1,0,2,0,1],
     [2,2,1,1,2,0,1,2,2,1,2,0,1,0,0,0,2,0,1,1,0,0,2,2,2,1,1,0,1,1,0],
     [0,2,0,1,0,2,1,1,2,1,0,1,2,2,0,0,0,1,0,2,0,0,2,2,2,0,1,0,1,2,1],
     [0,2,1,0,2,2,0,2,0,1,0,2,1,2,1,0,2,1,2,0,2,1,1,2,0,1,0,1,1,2,2],
     [1,1,0,1,2,1,2,0,2,2,1,0,0,0,2,1,2,0,0,2,0,2,1,1,1,1,2,2,0,1,0],
     [0,0,2,2,0,0,2,1,1,1,2,1,1,2,2,1,0,0,0,1,2,0,1,0,1,1,2,0,0,0,1],
     [2,0,1,2,2,2,0,1,2,0,0,1,1,2,1,0,0,2,2,1,1,1,0,2,2,0,1,1,0,0,1],
     [1,0,0,0,2,0,1,0,1,0,1,1,0,2,2,0,2,2,2,2,1,0,2,0,2,2,2,0,0,1,2],
     [2,1,0,2,1,1,1,1,0,0,1,1,1,2,0,1,2,2,2,0,1,1,2,2,0,1,0,0,2,0,0],
     [2,1,1,2,2,1,0,1,2,2,0,2,2,1,0,1,0,1,0,1,1,1,0,2,1,0,1,0,2,2,2],
     [0,1,1,0,1,1,2,0,0,2,1,1,1,1,2,0,1,1,1,0,2,2,1,2,0,1,0,2,0,1,1],
     [0,1,1,0,0,0,2,2,1,2,0,2,1,1,1,2,0,0,0,1,2,0,1,0,0,0,2,0,1,2,1],
     [2,0,1,0,0,1,1,0,1,0,2,0,2,0,1,2,2,1,1,1,0,0,0,0,2,2,0,1,0,0,2],
     [1,0,2,2,0,1,0,2,0,2,1,2,2,1,0,0,0,1,2,2,1,0,0,1,0,2,1,2,0,0,0],
     [1,2,2,1,2,0,1,1,0,1,2,0,2,0,0,2,1,0,1,2,0,1,1,2,0,2,1,1,2,0,0]]
  • $GR(4,K_{4},2)\geq 15$
    Adjacency matrix of a 4-edge-colored 14-vertex graph with each $K_{4}$ containing more than 2 colors.
    [[0,0,3,1,2,3,0,3,2,2,0,1,0,1],
     [0,0,2,0,1,1,1,3,0,3,2,2,3,1],
     [3,2,0,2,2,3,1,1,0,1,3,3,0,0],
     [1,0,2,0,0,2,2,3,1,3,0,3,1,2],
     [2,1,2,0,0,0,3,0,2,1,1,3,2,1],
     [3,1,3,2,0,0,1,0,1,2,2,0,1,3],
     [0,1,1,2,3,1,0,0,3,2,3,3,0,2],
     [3,3,1,3,0,0,0,0,2,2,1,2,2,1],
     [2,0,0,1,2,1,3,2,0,0,1,1,3,3],
     [2,3,1,3,1,2,2,2,0,0,3,1,1,0],
     [0,2,3,0,1,2,3,1,1,3,0,0,2,3],
     [1,2,3,3,3,0,3,2,1,1,0,0,0,1],
     [0,3,0,1,2,1,0,2,3,1,2,0,0,2],
     [1,1,0,2,1,3,2,1,3,0,3,1,2,0]]
  • $GR(4,K_{4},3)=10$
    Adjacency matrix of a 4-edge-colored 9-vertex graph with each $K_{4}$ containing more than 3 colors. There is only one such graph on 9 vertices.
    [[0,2,3,3,1,0,2,1,0],
     [2,0,1,0,3,3,2,0,1],
     [3,1,0,3,0,2,0,2,1],
     [3,0,3,0,2,1,1,0,2],
     [1,3,0,2,0,3,0,1,2],
     [0,3,2,1,3,0,1,2,0],
     [2,2,0,1,0,1,0,3,3],
     [1,0,2,0,1,2,3,0,3],
     [0,1,1,2,2,0,3,3,0]]